OpenStudy (anonymous):

A six-sided die is rolled 4 times in a row. What is the probability of getting 4 only on the last trial of this experiment? A. 1/36 B. 4/6 C. 1/1296 D. 125/1296

OpenStudy (anonymous):

B

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

1/6+1/6+1/6+1/6 ......... obviously 4/6

OpenStudy (anonymous):

ok each time u roll a die .. it has probibality of 4 1/6 ok

OpenStudy (anonymous):

so if u will roll it 4 times .. obviously every time u have probibality of 1/6 right .. so u just need to add all the probaibalities ,,.. that will give rise to 4/6

OpenStudy (anonymous):

ohhh ok i get it. thank you

OpenStudy (anonymous):

ur awesome at math lol :D

OpenStudy (anonymous):

may be not .. its simple u should someone else too

OpenStudy (anonymous):

so do i divide 75/1296 then for the final answer?

OpenStudy (kropot72):

Not really. Look at the answer choices.

OpenStudy (anonymous):

yeah i see em. But none of them say 75/1296

OpenStudy (anonymous):

well nevermind. I'll just take my best guess, thanks

OpenStudy (kropot72):

Sorry, my bad. The probability of not getting a four on a roll is 5/6. The probability of getting a four is 1/6. The probability of getting 4 only on the last trial of the experiment is: \[\frac{5}{6}\times\frac{5}{6}\times\frac{5}{6}\times\frac{1}{6}=\frac{125}{1296}\]

OpenStudy (anonymous):

thanks your the best

OpenStudy (anonymous):

helped me alot

OpenStudy (kropot72):

You're welcome :)